OpenStudy (anonymous):

Find the vertex of the parabola by completing the square. x^2-6x+8y=y

OpenStudy (anonymous):

Are you dure you wrote it right? Equation look like this y=2x^2-12x-25, is it just 8?

OpenStudy (anonymous):

its ok if he puts the y= last it doesnt really matter

OpenStudy (anonymous):

No the 8y? It has two variables x and y

OpenStudy (anonymous):

hmmmmmm

OpenStudy (anonymous):

If its like this and not 8y y=x^2-6x+8 y=(x^2 -6x+9) 8 -9 y=(x-3)^2 -1 Vertez = (3,-1)